OpenStudy (anonymous):

A bicycle manufacturing company makes a particular type of bike. Each child bike requires 4 hours to build and 4 hours to test. Each adult bike requires 6 hours to build and 4 hours to test. With the number of workers, the company is able to have up to 120 hours of building time and 100 hours of testing time for a week. If c represents child bikes and a represents adult bikes, determine which system of inequality best explains whether the company can build 10 child bikes and 12 adult bikes in the week.

OpenStudy (anonymous):

No, because the bike order does not meet the restrictions of 4c + 6a ≤ 120 and 4c + 4a ≤ 100 No, because the bike order does not meet the restrictions of 4c + 4a ≤ 120 and 6c + 4a ≤ 100 Yes, because the bike order meets the restrictions of 4c + 6a ≤ 120 and 4c + 4a ≤ 100 Yes, because the bike order meets the restrictions of 4c + 4a ≤ 120 and 6c + 4a ≤ 100

OpenStudy (anonymous):

I got it. The answer is C.

OpenStudy (texaschic101):

c = 10, a = 12 4c + 6a <= 120 4(10) + 6(12) <= 120 40 + 72 <= 120 112 <= 120 -- this is correct 4c + 4a <= 100 4(10) + 4(12) <= 100 40 + 48 <= 100 88 <= 100 -- this is correct I agree...the answer is C.

OpenStudy (ckaranja):

4c + 6a ≤ 120 4c + 4a ≤ 100 solving the inequality gives; 2a=20 a≤10 Replacing in equation 1; 4c+60≤120 4c≤60 c≤15 It can